The average height of a class students is 134.7 cm. The sum of all the heights is 3771.6 cm. There are 17 boys, how many girls?
Answer #1

Solution :

Mean = sum of scores/number of scores

134.7 = 3771.6/(x+17)

134.7(x+17) = 3771.6

134.7x + 2289.9 = 3771.6

134.7x = 3771.6 - 2289.9

  134.7x = 1481.7

x = 1481.7/134.7

x = 11